Lead metal can be extracted from a mineral called galena, which contains 86.6%% lead by mass. A particular ore contains 68.5%% galena by mass.

Part A

If the lead can be extracted with 92.5%% efficiency, what mass of ore is required to make a lead sphere with a 7.00 cmcm radius?

Express your answer with the appropriate units.
Answer #1

lead sphere with a 7.00 cm radius :

volume of such sphere , V = 4/3 πr³ = 4/3 π*(7 cm)³ = 1436.8 cm3.

density of Pb : 11.34 g/cm3 .

mass of Pb : 1436.8 cm3 * 11.34 g/cm3 = 16.3 kg.

If the lead can be extracted with 92.5%% efficiency :

then mass of lead in ore required : 16.3 kg/0.925 = 17.6 kg

Lead metal can be extracted from a mineral called galena, which contains 86.6%% lead by mass.

mass of mineral (PbS) : 17.6 kg/0.866 = 20.34 Kg

A particular ore contains 68.5% galena by mass :

thus , mass of ore is required : 20.34 Kg/0.685 = 29.7 Kg.